# TRELLISWorld: Training-Free World Generation from Object Generators

### v1

@@ 1 Introduction
Recent advances in 3D content generation, ranging from SDS-based methods distilling 2D priors into neural fields (Poole et al., 2022; Wang et al., 2023; 2022; Tang et al., 2024b), to multi-view diffusion for geometric consistency (Liu et al., 2023b; Shi et al., 2023; 2024; Liu et al., 2024), and emerging 3D-native models like DiTs (Peebles & Xie, 2023; Hong et al., 2024b; Lai et al., 2025; Zhang et al., 2024b), have significantly improved 3D object synthesis. However, these methods are predominantly limited to single-object generation, with limited progress toward generating entire 3D scenes.
@@
Recently, SynCity (Engstler et al., 2025) demonstrates promising results in adapting object generators for city-scale scene generation. However, its dependence on 2D inpainting (Lugmayr et al., 2022) introduces a fundamental limitation. Errors in the image domain can propagate and destabilize the 3D reconstruction. This raises a natural question: can we instead generate scenes directly in 3D space, bypassing the need for 2D intermediate representations?
@@
  • Training-free and editable: Taking advantages of 3D environments’ multi-scale signal structure, our approach requires no scene-level dataset or fine-tuning. It inherits editability and generalization capabilities from the underlying object-level generator, e.g., TRELLIS (Xiang et al., 2025).
@@
  • Scalable and smooth: Compared to prior methods, our tile-wise approach is computationally efficient, blends overlapping regions smoothly, and enables generations of significantly larger and more coherent scenes.
@@ 2 Related Work
Figure 2: Illustration of our tiled diffusion process. We first split the scene noise into multiple tiles to denoise each tile in parallel. Then we take the weighted average described in Equation 1 for each tile and aggregate the result to obtain the scene noise for previous timesteps. This process is detailed by Equation 2.
@@ 2.1 Foundation of Reconstruction and Object Generation
Benefiting from recent advances in 3D representations for reconstruction, such as NeRF (Mildenhall et al., 2020; Yu et al., 2021; Müller et al., 2022; Shue et al., 2022; Chen et al., 2022; Kerbl et al., 2023; Zhang et al., 2020; Barron et al., 2021; 2022; 2023), Score Distillation Sampling (SDS)-based methods (Poole et al., 2022; Wang et al., 2023; 2022; Tang et al., 2024b) distill the knowledge of 2D diffusion models into the creation of 3D objects using differentiable rendering. Specifically, NeRF++ (Zhang et al., 2020), combined with ProlificDreamer (Wang et al., 2023), demonstrates the first possibility of generating 3D scenes by distilling 2D diffusion with an added density prior. Later, multi-view diffusion methods (Liu et al., 2023b; Shi et al., 2023; 2024; Liu et al., 2024) were proposed to address the Janus (multi-face) problem. However, as SDS depends on test-time optimization, object generation can take up to 40 minutes. To overcome this, methods like LRM have been proposed to generate 3D representations directly from images (Liu et al., 2023a; Hong et al., 2024b; Gao et al., 2024; Lai et al., 2025) and texts (Jun & Nichol, 2023; Hong et al., 2024a; Zhang et al., 2024b). Although none of these works aim to generate large-scale scenes, their technologies lay the foundation for 3D scene generation.
@@ 2.2 Scene Generation Based on 2D Generation
While SDS-based methods above often focus on novel view synthesis or scene reconstruction, another line of work leverages the capabilities of 2D diffusion models without relying on test-time optimization. Early works extend a single image autoregressively into a video sequence, guided by camera trajectories (Liu et al., 2021; Li et al., 2022; Chai et al., 2023; Cai et al., 2023). These methods often project depth-estimated results into point clouds and fill in the missing regions under different camera extrinsics. While initial work focused mainly on natural scenes, the introduction of generic 2D diffusion models expanded the domain (Fridman et al., 2023; Shriram et al., 2025; Chung et al., 2023), and the integration of large language models (LLMs) enabled more diverse scene generation (Yu et al., 2024; 2025; Team et al., 2025). However, these works are limited to generating panoramic images or incomplete 3D representations that can only be viewed from a restricted set of camera extrinsics. A generic 3D generator capable of producing complete meshes remains necessary.
@@ 2.3 Scene Generation Based on 3D Scene-Native Generation
3D-native generation methods are promising alternatives to produce complete 3D representations, but they often rely on domain-specific training and are unable to generate generic scenes using natural language prompts (Lee et al., 2024; Tang et al., 2024a; Meng et al., 2025; Lee et al., 2025). For instance, works like InfiniCity (Lin et al., 2023; Xie et al., 2025b; a) depend heavily on RGB-D, semantic, or normal maps derived from satellite imagery. CityDreamer4D (Xie et al., 2025a) decomposes the generation task into multiple sub-tasks: layout, background, buildings, vehicles, and roads, each handled by a separate neural network. BlockFusion (Wu et al., 2024) uses autoregressive inference, while MIDI (Huang et al., 2025) employs multi-instance attention to scale object-level generation to scenes, both trained on dedicated 3D datasets (Fu et al., 2021; 2020). In contrast to publicly accessible 3D object datasets like Objaverse (Deitke et al., 2022; 2023), which contain over 10 million internet-sourced objects, the largest 3D scene dataset, FurniScene, contains only 100k rooms with 89 object classes (Zhang et al., 2024a). To the best of our knowledge, no openly available generic 3D scene dataset currently exists. Thus, a method that does not rely on curated 3D scene datasets but can still generate 3D representations across domains is highly desirable.
@@ 2.4 Scene Generation Based on 3D Object Generation
There are two categories of works that utilize 3D object generators to build scenes. The first category focuses on generating individual 3D objects and uses LLMs, visual-language models (VLMs), or image-based techniques to infer plausible object positions and orientations (Feng et al., 2023; Wu et al., 2025; Li et al., 2025). GALA3D (Zhou et al., 2024) employs SDS and additional physical losses to refine LLM-generated scene layouts from textual descriptions. CAST (Yao et al., 2025) constructs a relational graph, constraint graph, and multiple masks from a single image to generate both object instances without occlusion and their spatial arrangement, using physical losses for consistency. However, these methods either depend on image inputs, limiting composition outside the view frustum, or rely on LLM reasoning, which often fails to produce accurate coordinates and complex inter-object relationships.
The second category relies on the inpainting capability of 3D diffusion models to compose and blend scenes from object-level generations. Most relevant to our work, SynCity (Engstler et al., 2025) generates object tiles autoregressively through a loop of 2D inpainting, 3D generation, and rendering. After generation, it fixes seams between tiles using 3D inpainting. To ensure consistency between 3D tiles and to avoid occlusions during 2D inpainting, multiple heuristics, such as cutting off parts of generated meshes for occlusion-free renderings, are applied. However, these heuristics reduce generalizability beyond urban scenes and increase failure cases.
@@ 3 Method
The core of TRELLISWorld is a tiled diffusion with cosine blending. We first formulate the problem in subsection 3.1, then give a general method in subsection 3.2. We then describe how we actually implement this method using TRELLIS in subsection 3.3.
@@ 3.1 Problem Formulation
Given a text-conditioned 3D generative diffusion model $\theta$ (a velocity field (Lipman et al., 2023)) that is capable of generating a 3D structure of size $S^{3}$ from a text prompt $p$, our goal is to generate a large-scale 3D world with arbitrary size $(X\times Y\times Z)>>S^{3}$ that is consistent with the prompt. To simplify our explanation, without loss of generalizability, we assume $\theta$ is a pixel-diffusion, meaning that the forward and reverse diffusion process are on actual value instead of on compressed latent by autoencoders. Therefore, each object sample can be represented with a tensor $\mathbb{R}^{S^{3}}$. We further assume that $\theta$ is trained on general object distribution. To generate worlds using object-level 3D generator, we leverage the local generation capability of $\theta$ while ensuring global coherence through careful conditioning and blending techniques.
@@ 3.2 Tiled Diffusion
We first initialize the entire world $W$ of size $(X,Y,Z)$ with Gaussian noise $W\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. We then divide the world into overlapping cubic tiles $\{w_{i}\}$ of size $(S,S,S)$ with a stride of $(s,s,s)$ where $s<S$ to ensure overlap between adjacent tiles. For each diffusion step, we process each tile $\{w_{i}\}$ in parallel and then aggregate the weighted average result to update the world $W$. The weight for each tile is defined by a 3D cosine mask that emphasizes the center of the tile and tapers off towards the edges, ensuring smooth transitions between adjacent tiles. See Figure 2 for visualization.
Formally, let $v^{(x,y,z)}$ to be a voxel that has global position $(x,y,z)$ and let $f_{w_{i}}:\mathbb{N}^{3}\to\mathbb{Z}^{3}$ be a function to map global position into local position relative to tile $w_{i}$. Intuitively, if the resulting position $f_{w_{i}}(\cdot)$ is in the set $\{0,...,S-1\}^{3}$, then the tile $w_{i}$ covers the voxel at this input global position.
@@
Then, the update rule for each voxel $v^{(x,y,z)}$ in the world $W$ at diffusion step $t$ is given by:
@@ 3.3 Implementation
We build our method based on TRELLIS-text (Xiang et al., 2025), which is a text-conditioned diffusion transformer for 3D object generation. To perform object-level inference with TRELLIS, the process begins by denoising a $16^{3}$ noised latent using the TRELLIS structure diffusion transformer $\theta_{1}$. The resulting latent is decoded by TRELLIS sparse structure decoder to a $64^{3}$ occupancy grid (SS). This dense $64^{3}$ tensor is converted into a noisy sparse tensor by retaining only regions where the occupancy value exceeds zero and apply noise $\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ onto those regions. The sparse tensor is subsequently denoised using the TRELLIS structure latent diffusion transformer $\theta_{2}$, resulting in a structured latent representation (SLAT). Finally, the SLAT is decoded using the TRELLIS structure latent Gaussian decoder to produce a Gaussian Splatting representation (Kerbl et al., 2023).
As described above, TRELLIS is a multi-stage latent diffusion model where $\theta_{1}$ operates on dense tensors and $\theta_{2}$ operates on sparse tensors. We perform tiled diffusion with minimal modification on both stages: the input tiles to $\theta_{1},\theta_{2}$ are in encoded latent space and the corresponding masks are down-sampled by $4\times$. The blending and aggregation steps described in subsection 3.2 are performed in the latent space, followed by decoding back to the voxel space and/or Gaussian Splatting space after the diffusion process is complete. Importantly, such decoding should also be done in a tiled manner. Since TRELLIS structure latent Gaussian decoder is not a probabilistic model, we set the stride $s=S$ to disable blending.
@@ 4 Experiments
We ablate our method in subsection 4.1 and compare blend quality, perceptual alignments, and computational cost with SynCity in subsection 4.2 and subsection 4.3. All experiments are conducted using classification-free guidance (Ho & Salimans, 2022) $cfg=7.5$, stride $s=\frac{S}{2}$ and diffusion step size $25$ on Euler sampler.
@@ 4.1 Ablation
A naive approach is to generate the world by stitching together multiple object generation results autoregressively, ensuring context alignment of neighboring tiles using inpainting. However, this leads to less coherent generation at the tile edges, as shown in Figure 3.
@@
Figure 4: Comparison (not cherry-picked) showing the effectiveness of blending. (a) Without blending, the ”room” example tends to generate walls around tile borders, and the ”lego tile” example produces a colored edge along tile borders, which is undesirable. (b) Tile borders become less noticeable with blending.
@@
Tiled Decoder
Figure 5: Comparison example with 3x2x1 city tiles for (a) decoding using our tiled decoding method and (b) decoding the entire generation at once. We observe severe artifacts when decoding without our tiled decoder.
We compare decoding using our tiled decoder with decoding the entire world at once in Figure 5. Removing the tiled decoder shows artifacts on generated Gaussian Splatting.
@@ 4.2 Qualitative Comparison
Figure 6: Qualitative comparison (not cherry-picked) between SynCity (Engstler et al., 2025) and TRELLISWorld (our method). All generations use 4x3x1 tiles under the same Gaussian Splatting resolution. Our method demonstrates seamless blending between tiles, whereas tile boundaries in SynCity are easily noticeable.
We compare our method with SynCity (Engstler et al., 2025). Figure 6 showcases generation results across multiple prompts using the same 4x3x1 tile size. Our method tends to generate larger scenes and provides more natural blending across tiles compared to SynCity. For additional generation results of TRELLISWorld, see Figure 11 and Figure 12 in Appendix A.
Figure 7: Limitations of world generation using autoregressive image inpainting methods described by SynCity (Engstler et al., 2025). In SynCity, to prevent tall buildings from previous tiles from occluding the next tile, the previously generated result in (a) is cropped to (b), producing an inpainted image with artifacts in (c). The object generator is then conditioned on image (c) to generate Gaussian Splatting (d), which contains inherited artifacts.
Furthermore, our method is more robust compared to autoregressive methods based on image inpainting. For example, SynCity relies on heuristics such as cropping the 3D generation to avoid occlusion before applying 2D inpainting. Such heuristics are prone to failure when the 2D diffusion model mimics the cropped content, producing 3D tiles with artifacts, as shown in Figure 7.

### v2

@@ 1 Introduction
Recent advances in 3D content generation, ranging from SDS-based methods distilling 2D priors into neural fields (sds; vsd; sjc; dreamgaussian), to multi-view diffusion for geometric consistency (z123; z123pp; mvdream; syncdreamer), and emerging 3D-native models like DiTs (dit; lrm; hunyuan3d; clay), have significantly improved 3D object synthesis. However, these methods are predominantly limited to single-object generation, with limited progress toward generating entire 3D scenes.
@@
Recently, SynCity (syncity) demonstrates promising results in adapting object generators for city-scale scene generation. However, its dependence on 2D inpainting (inpaint) introduces a fundamental limitation. Errors in the image domain can propagate and destabilize the 3D reconstruction. This raises a natural question: can we instead generate scenes directly in 3D space, bypassing the need for 2D intermediate representations?
@@
  • Training-free and editable: Taking advantage of 3D environments’ multi-scale signal structure, our approach requires no scene-level dataset or fine-tuning. It inherits editability and generalization capabilities from the underlying object-level generator, e.g., TRELLIS (trellis).
@@
  • Scalable and smooth: Compared to prior methods, our tile-wise approach is computationally efficient, blends overlapping regions smoothly, and enables the generation of significantly larger and more coherent scenes.
@@ 2.1 Foundation of Reconstruction and Object Generation
Benefiting from recent advances in 3D representations for reconstruction, such as NeRF (nerf; plenoxel; ngp; triplane; tensorf; gsplat; nerfpp; mipnerf; mipnerf360; zipnerf), Score Distillation Sampling (SDS)-based methods (sds; vsd; sjc; dreamgaussian) distill the knowledge of 2D diffusion models into the creation of 3D objects using differentiable rendering. Specifically, NeRF++ (nerfpp), combined with ProlificDreamer (vsd), demonstrates the first possibility of generating 3D scenes by distilling 2D diffusion with an added density prior. Later, multi-view diffusion methods (z123; z123pp; mvdream; syncdreamer) were proposed to address the Janus (multi-face) problem. However, as SDS depends on test-time optimization, object generation can take up to 40 minutes. To overcome this, methods like LRM have been proposed to generate 3D representations directly from images (12345; lrm; cat3d; hunyuan3d) and text (shape-e; 3dtopia; clay). Although none of these works aim to generate large-scale scenes, their technologies lay the foundation for 3D scene generation.
@@ 2.2 Scene Generation Based on 2D Generation
While SDS-based methods above often focus on novel view synthesis or scene reconstruction, another line of work leverages the capabilities of 2D diffusion models without relying on test-time optimization. Early works extend a single image autoregressively into a video sequence, guided by camera trajectories (infinite-nature; infinite-nature-zero; persist-nature; diffdreamer). These methods often project depth-estimated results into point clouds and fill in the missing regions under different camera extrinsics. While initial work focused mainly on natural scenes, the introduction of generic 2D diffusion models expanded the domain (scenescape; realmdreamer; luciddreamer-c), and the integration of large language models (LLMs) enabled more diverse scene generation (wonderjourney; wonderworld; hunyuanworld). However, these works are limited to generating panoramic images or incomplete 3D representations that can only be viewed from a restricted set of camera extrinsics. A generic 3D generator capable of producing complete meshes remains necessary.
@@ 2.3 Scene Generation Based on 3D Scene-Native Generation
3D-native generation methods are promising alternatives to produce complete 3D representations, but they often rely on domain-specific training and are unable to generate generic scenes using natural language prompts (semcity; diffuscene; scenegen; nuiscene). For instance, works like InfiniCity (infinicity; gaussiancity; citydreamer4d) depend heavily on RGB-D, semantic, or normal maps derived from satellite imagery. CityDreamer4D (citydreamer4d) decomposes the generation task into multiple sub-tasks: layout, background, buildings, vehicles, and roads, each handled by a separate neural network. BlockFusion (blockfusion) uses autoregressive inference, while MIDI (midi) employs multi-instance attention to scale object-level generation to scenes, both trained on dedicated 3D datasets (3dfront; 3dfuture). In contrast to publicly accessible 3D object datasets like Objaverse (objaverse; objaverse-xl), which contain over 10 million internet-sourced objects, the largest 3D scene dataset, FurniScene, contains only 100k rooms with 89 object classes (furniscene). To the best of our knowledge, no openly available generic 3D scene dataset currently exists. Thus, a method that does not rely on curated 3D scene datasets but can still generate 3D representations across domains is highly desirable.
@@ 2.4 Scene Generation Based on 3D Object Generation
There are two categories of works that utilize 3D object generators to build scenes. The first category focuses on generating individual 3D objects and uses LLMs, visual-language models (VLMs), or image-based techniques to infer plausible object positions and orientations (layoutgpt; diorama; phip-g). GALA3D (gala3d) employs SDS and additional physical losses to refine LLM-generated scene layouts from textual descriptions. CAST (cast) constructs a relational graph, a constraint graph, and multiple masks from a single image to generate both object instances without occlusion and their spatial arrangement, using physical losses for consistency. However, these methods either depend on image inputs, limiting composition outside the view frustum, or rely on LLM reasoning, which often fails to produce accurate coordinates and complex inter-object relationships.
The second category relies on the inpainting capability of 3D diffusion models to compose and blend scenes from object-level generations. Most relevant to our work, SynCity (syncity) generates object chunks autoregressively through a loop of 2D inpainting, 3D generation, and rendering. After generation, it fixes seams between chunks using 3D inpainting. To ensure consistency between 3D chunks and to avoid occlusions during 2D inpainting, multiple heuristics, such as cutting off parts of generated meshes for occlusion-free renderings, are applied. However, these heuristics reduce generalizability beyond urban scenes and increase failure cases.
@@ 3 Method
Figure 2: Illustration of our tiled diffusion process. We first split the scene noise into multiple tiles to denoise each tile in parallel. Then we take the weighted average described in Equation 1 for each tile and aggregate the result to obtain the scene noise for previous timesteps. This process is detailed in Equation 2.
The core of TRELLISWorld is a tiled diffusion with cosine blending. We first formulate the problem in subsection 3.1, then give a general method in subsection 3.2. We then describe how we implement this method using TRELLIS in subsection 3.3.
@@ 3.1 Problem Formulation
Given a text-conditioned 3D generative diffusion model $\theta$ (a velocity field (flow-matching)) that is capable of generating a 3D structure of size $S^{3}$ from a text prompt $p$, our goal is to generate a large-scale 3D world of arbitrary size $(X\times Y\times Z)\gg S^{3}$ that is consistent with the prompt. To simplify our explanation, without loss of generality, we assume $\theta$ is a pixel-diffusion model, meaning that the forward and reverse diffusion processes operate on the actual values instead of on compressed latents produced by autoencoders. Therefore, each object sample can be represented with a tensor in $\mathbb{R}^{S^{3}}$. We further assume that $\theta$ is trained on a general object distribution. To generate worlds using an object-level 3D generator, we leverage the local generation capability of $\theta$ while ensuring global coherence through careful conditioning and blending techniques.
@@ 3.2 Tiled Diffusion
We first initialize the entire world $W$ of size $(X,Y,Z)$ with Gaussian noise $W\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. We then divide the world into overlapping cubic tiles $\{w_{i}\}$ of size $(S,S,S)$ with a stride of $(s,s,s)$, where $s<S$ to ensure overlap between adjacent tiles. For a fixed scene, different settings of $s$ give different numbers of tiles. Thus, we use the word ”chunk” to denote the absolute size of the scene. For each diffusion step, we process each tile $\{w_{i}\}$ in parallel and then aggregate the weighted average result to update the world $W$. The weight for each tile is defined by a 3D cosine mask that emphasizes the center of the tile and tapers off towards the edges, ensuring smooth transitions between adjacent tiles. See Figure 2 for visualization.
Formally, let $v^{(x,y,z)}$ be a voxel that has global position $(x,y,z)$ and let $f_{w_{i}}:\mathbb{N}^{3}\to\mathbb{Z}^{3}$ be a function mapping global positions into local positions relative to tile $w_{i}$. Intuitively, if the resulting position $f_{w_{i}}(\cdot)$ lies in the set $\{0,...,S-1\}^{3}$, then the tile $w_{i}$ covers the voxel at this input global position.
@@
Then, the update rule for each voxel $v^{(x,y,z)}$ in the world $W$ at diffusion step $t$ is given by the formula:
@@
where $\mathcal{O}(\Delta t^{2})$ denotes the Big-O error resulting from Euler discretization, and the subscript $\left[\cdot\right]_{f_{w_{i}^{(t)}}(x,y,z)}$ denotes that $\theta$ operates in the local tile space.
@@ 3.3 Implementation
We build our method based on TRELLIS-text (trellis), which is a text-conditioned diffusion transformer for 3D object generation. To perform object-level inference with TRELLIS, the process begins by denoising a $16^{3}$ noised latent using the TRELLIS structure diffusion transformer $\theta_{1}$. The resulting latent is decoded by TRELLIS sparse structure decoder to a $64^{3}$ occupancy grid (SS). This dense $64^{3}$ tensor is converted into a noisy sparse tensor by retaining only regions where the occupancy value exceeds zero. The noise $\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ is applied onto those regions. The sparse tensor is subsequently denoised using the TRELLIS structure latent diffusion transformer $\theta_{2}$, resulting in a structured latent representation (SLAT). Finally, the SLAT is decoded using the TRELLIS structure latent Gaussian decoder to produce a Gaussian Splatting representation (gsplat).
Tiled Diffusion
As described above, TRELLIS is a multi-stage latent diffusion model where $\theta_{1}$ operates on dense tensors and $\theta_{2}$ operates on sparse tensors. We perform tiled diffusion with minimal modification on both stages: the input tiles to $\theta_{1},\theta_{2}$ are in encoded latent space and the corresponding masks are down-sampled by $4\times$. The blending and aggregation steps described in subsection 3.2 are performed in the latent space.
Tiled Decoder
The resulting latent is decoded back to the voxel space and/or Gaussian Splatting space after the diffusion process is complete. Importantly, such decoding should also be done in a tiled manner. Since TRELLIS structure latent Gaussian decoder is not a probabilistic model, we set the stride $s=S$ to disable blending.
@@ 4 Experiments
We ablate our method in subsection 4.1 and compare blend quality, perceptual alignments, and computational cost with SynCity in subsection 4.2 and subsection 4.3. Unless otherwise noted, all experiments are conducted using classification-free guidance (cfg) $cfg=7.5$, stride $s=\frac{S}{2}$ and diffusion step size $25$ on the Euler sampler.
@@ 4.1 Ablation
(a) Autoregressive
Refer to caption Refer to caption Refer to caption
(b) TRELLISWorld
Refer to caption Refer to caption Refer to caption
Desert Cyberpubk Forest
@@
A naive approach is to generate the world by stitching together multiple object generation results autoregressively, ensuring context alignment of neighboring chunks using inpainting. However, this leads to less coherent generation at the chunk edges, as shown in Figure 3.
Tiled Decoder
(a) with tiled decoder
(b) without tiled decoder
Figure 4: Comparison example with 3x2x1 city chunks for (a) decoding using our tiled decoding method and (b) decoding the entire generation at once. We observe severe artifacts when decoding without our tiled decoder.
We compare decoding using our tiled decoder with decoding the entire world at once in Figure 4. Removing the tiled decoder shows artifacts in the generated Gaussian Splatting.
@@
(a) Average Blending
Refer to caption Refer to caption
(b) TRELLISWorld
Refer to caption Refer to caption
Room Lego Tile
Figure 5: Comparison (not cherry-picked) showing the effectiveness of blending. (a) With average blending, the ”room” example tends to generate walls around tile borders, and the ”lego tile” example produces a colored edge along tile borders, which is undesirable. (b) Tile borders become less noticeable with blending.
@@ 4.2 Qualitative Comparison
(a) SynCity
Refer to caption Refer to caption Refer to caption
(b) TRELLISWorld
Refer to caption Refer to caption Refer to caption
Forest Cyberpubk Minecraft
Figure 6: Qualitative comparison (not cherry-picked) between (a) SynCity (syncity) and (b) TRELLISWorld (our method). All generations use a 4x3x1 chunk layout under the same Gaussian Splatting resolution. Our method demonstrates seamless blending between tiles, whereas chunk boundaries in SynCity are easily noticeable.
We compare our method with SynCity (syncity). Figure 6 showcases generation results across multiple prompts using the same 4x3x1 chunk layout. Our method tends to generate larger scenes and provides more natural blending across tiles compared to SynCity. For additional generation results of TRELLISWorld, see Figure 18 and Figure 19 in Appendix A.
(a) previous tiles
(b) apply cropping
(b) inpainted image
(b) generated next tile
Figure 7: Limitations of world generation using autoregressive image inpainting methods described by SynCity (syncity). In SynCity, to prevent tall buildings from previous chunks from occluding the next chunk, the previously generated result in (a) is cropped to (b), producing an inpainted image with artifacts in (c). The object generator is then conditioned on image (c) to generate Gaussian Splatting (d), which contains inherited artifacts.
Furthermore, our method is more robust compared to autoregressive methods based on image inpainting. For example, SynCity relies on heuristics such as cropping the 3D generation to avoid occlusion before applying 2D inpainting. Such heuristics are prone to failure when the 2D diffusion model mimics the cropped content, producing 3D chunks with artifacts, as shown in Figure 7.